# Structured Packing in LLM Training Improves Long Context Utilization

### v8

@@ C.1 SPLiCe parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 16 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters for $270$M models. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
@@
We also evaluated the influence of BOS and EOS tokens on the performance of trained models. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively slightly decrease the number of separating tokens compared to the Baseline. However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
@@ C.2 Importance of separating tokens and in-domain sampling
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training examples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens and concatenation of documents from the same domain, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the C dataset (sampling without replacement). The results are shown in Table 17. We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains. When concatenating the documents, we concatenate them within the domain (C language in this case).
Table 17: Perplexity evaluation of two methods of organizing the data. Baseline – example packing, document equals training example. DomRnd – concatenate documents within the domain into examples of length bounded by $120$K characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in DomRnd does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation. The training data is a 50/50 mixture of RedPajama, organized in a standard way, and C data, organized using a method of choice.
@@
C DomRnd 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
@@ Appendix D Key-Value retrieval task
Figure 2 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in [Liu et al., 2023]. This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
Extract the value corresponding to the specified key in the JSON object below. JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
We noted that FoT-trained models struggle with this task. This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits. Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3. For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 3.
Figure 3: Performance on a smaller version of key-value retrieval task from [Liu et al., 2023]. We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama [Rozière et al., 2023] method of context extension. Each position was evaluated using $500$ examples.
@@ Appendix E Perplexity improvements
In Figure 4 we present perplexity improvements of $3$B FoT SPLiCe BM25 over Baseline. Figure 5 shows the evolution of SPLiCe model perplexity during the training. We follow [Anthropic, 2023] and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 5: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 4 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Shuffling
Table 18: Average classification performance on TREC [Li and Roth, 2002, Hovy et al., 2001]. We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
TREC
Model
Context length
(# of examples)
SPLiCe-no-shuf SPLiCe-shuf
$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
Appendix G Data preparation
G.1 Evaluation data
@@
G.2 Train data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$K files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used to organize the data. For C, we have performed a run on three packs and provided results and standard deviation in Table 5. For large models, we run the methods on several packs and concatenate the results into a single dataset. For natural language datasets, we extract a random subset of documents.
Appendix H Faiss parameters
Our experiments with SPLiCe Cont utilize Faiss [Johnson et al., 2019] for fast approximate inner-product search. To be more precise, we use the "IVF8192,Flat" index that we train on 262144 examples coming from the dataset.
Appendix I Detailed accuracy improvements
@@
$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$. Below we report the histograms of $\delta(c)$. In Table 1 and Table 2 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
@@
Figure 6: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
@@
Figure 7: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on DBPedia. We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
Appendix J Longer context
In Table 20, we present the perplexity results for $270$M parameter models trained as described in Section 3.3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
@@
Altered Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
SPLiCe Cont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
SPLiCe Repo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
Baseline 5.01 2.79 2.82 2.51 2.94 3.07
Appendix K Data distribution property
Chan et al. [2022], Han et al. [2023] shows that the "burstiness" property of training data benefits a model’s generalisation. Specifically, "burstiness" presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency. We also observe a burstiness property of data in SPLiCe, as shown in Table 21.
Table 21: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
Training Data Context Length Method Zipf’s Coefficient
C 2K SPLiCe BM25 ${1.667}_{(0.139)}$
Baseline ${1.717}_{(0.152)}$
StackEx 2K SPLiCe BM25 ${1.836}_{(0.088)}$
Baseline ${1.865}_{(0.074)}$
C 32K SPLiCe BM25 ${1.512}_{(0.055)}$
Baseline ${1.593}_{(0.025)}$
StackEx 32K SPLiCe BM25 ${1.643}_{(0.026)}$
Baseline ${1.664}_{(0.013)}$
@@
Table 22: We additionally evaluate our $7$B and $13$B parameter models on HumanEval [Chen et al., 2021]. For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance. For $13$B, both SPLiCe and Baseline increase the performance. We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion. Instead, evaluate all models in the same way. Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in [Rozière et al., 2023].
@@
Baseline 26.8

### v9

@@ C.1 SPLiCe Parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 23 indicates that $k=1$ is the best choice for naturally ocurring long documents whereas Table 24 shows that in case of RAG-style input greater values of $k$ achieve better performance, though the differences are rather small. We found that changing the order of documents in training samples hardly matters for $270$M models. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
@@
Method Code Code &
arXiv C RAG-style Python Code arXiv
SPLiCeBM25 k=1 5.46 3.20 2.81 2.94 3.10
SPLiCeBM25 k=16 5.49 3.18 2.85 2.98 3.14
Table 24: Ablation of SPLiCe hyper-parameter $k$, we can observe that increasing $k$ can bring benefits on RAG-style input data (input created by retrieving related documents with SPLiCe $k=16$) at a cost of performance on natural long documents. Models were tuned on a 50/50 mixture of C prepared with SPLiCe and RedPajama
We also evaluated the influence of BOS and EOS tokens on the performance of trained models. As both Repo and SPLiCe methods concatenate documents to create training samples, they effectively slightly decrease the number of separating tokens compared to the EP. However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
@@ C.2 Importance of Separating Tokens and In-Domain Sampling
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups, training samples are separated by BOS and EOS tokens. As SPLiCe methods concatenate documents to create training samples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens and concatenation of documents from the same domain, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the C dataset (sampling without replacement). The results are shown in Table 25. We note that the difference between the EP and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains. When concatenating the documents, we concatenate them within the domain (C language in this case).
@@
C WithinDomEP 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
EP 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
Table 25: Perplexity evaluation of two methods of organizing the data. EP – example packing, document equals training sample. WithinDomEP – concatenate documents within the domain into examples of length bounded by $120$K characters. Training samples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in WithinDomEP does not help the model to utilize extended context. Experiments were performed using the $270$M parameter model. We performed three runs on different subsets of C to provide mean and standard deviation. The training data is a 50/50 mixture of RedPajama, organized in a standard way, and C data, organized using a method of choice.
@@ Appendix D Key-Value Retrieval Task
Figure 3 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in (Liu et al. 2023). This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
Extract␣the␣value␣corresponding␣to␣the␣specified␣key␣in␣the␣JSON␣object␣below.
JSON␣data:
{"4ef217b7-6bc0-48c6-af35-2765f1e730f3":␣"068192b7-16b1-40e0-8495-61c63f979d50",
␣"cd6b8bdc-bc6c-4490-acb4-bc187a2dccba":␣"7364a26e-289f-4968-93d3-b273e882bdee",
␣"7d057372-4ab8-4811-8110-658c3f19fff4":␣"3ad075c5-b567-4201-85a7-cb31a0c91540",
␣"c62e192d-45e6-4646-bb88-1529c73256c9":␣"f0411644-1f6d-42a6-8af8-f06da66efc77",
␣"06134e93-e158-490e-a66c-8e3b98e12735":␣"50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",
␣"3286f978-4270-4b54-8bfa-540d7e0772e6":␣"075cc716-1836-4f90-9be3-53e3d4ec6585",
␣"4701aa05-c523-4b89-9700-64ab9c37c537":␣"49d86354-74c4-4256-9b3a-35e6e2b80d00",
␣"c8895805-e574-4f13-9fe5-89da1d8c4748":␣"cc91af7f-8509-4bdc-bad7-2646af68e6d2"}
␣"4701aa05-c523-4b89-9700-64ab9c37c537":
We noted that FoT-trained models struggle with this task. This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits. Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 4, and 4. For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 4.
Figure 4: Performance on a smaller version of key-value retrieval task from (Liu et al. 2023). We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context. For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al. 2023) method of context extension. Each position was evaluated using $500$ examples.
@@ Appendix E Perplexity Improvements
In Figure 5 we present perplexity improvements of $3$B FoT SPLiCeBM25 over EP. Figure 6 shows the evolution of SPLiCe model perplexity during the training. We follow (Anthropic 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
Figure 5: Perplexity improvement with SPLiCe against the EP of the final models (after $21$k training steps). We bucket tokens by their positions in the document and calculate the average. Each dot is the difference of the averages of the SPLiCe and EP models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 6: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 5 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training. For more see Table 26.
Steps 4K 8K 16K
TREC 64.2 $\pm$ 5.3 72.9 $\pm$ 4.1 72.9 $\pm$ 4.1
Table 26: Evaluation of the 3B SPLiCe model on 780-shot TREC using 10 seeds and different checkpoints. See Figure 6 for perlexity.
Appendix F Data Preparation
F.1 Evaluation Data
@@
F.2 Train Data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section F.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$K files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used to organize the data. For C, we have performed a run on three packs and provided results and standard deviation in Table 7. For large models, we run the methods on several packs and concatenate the results into a single dataset. For natural language datasets, we extract a random subset of documents.
Appendix G Faiss Parameters
Our experiments with SPLiCeCont utilize Faiss (Johnson, Douze, and Jégou 2019) for fast approximate inner-product search. To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
Appendix H Detailed Accuracy Improvements
@@
$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{EP}}(c),$
where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{EP}}(c)$ are the accuracies of the model trained with SPLiCe and EP respectively on the random choice of in-context examples $c$. Below we report the histograms of $\delta(c)$. In Table 1 and Table 2 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
@@
Figure 7: Histograms of accuracy improvement of SPLiCeBM25 over EP on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (EP) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the EP.
@@
Figure 8: Histograms of accuracy improvement of SPLiCeBM25 over EP on DBPedia. We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
Appendix I Results of $3$B Models on MMLU and GSM8K
We present the missing results in Table 28.
Model MMLU GSM8K
STEM HUM All
$3$B ST CHKPT 24.4 25.6 25.9 4.3
@@ Appendix I Results of $3$B Models on MMLU and GSM8K
[…]
Table 28: We evaluate our models on MMLU ($5$-shot), GSM8K ($8$-shot CoT). We provide an additional comparison with their starting checkpoint. In the main paper we note that the results of $3$B parameter model starting checkpoint are close to random and were moved this Appendix. See Table 3 for results regarding larger models.
Appendix J Ordering of Examples
We typically use the identity ordering as Order in Algorithm 1 to merge documents into a single context, as we found that in most cases, it performs best. We found that random shuffling is slightly better in some cases. Specifically, this is the case of large $7$B CodeLlama models, see Table Table 29. We hypothesize that random ordering forces the model to make use of the full space of the RoPe positional encoding. Whereas the identity ordering, also used in (Shi et al. 2024), skews the model toward paying more attention to fragments of text that are not too far away. This suggests an interesting research direction on the intersection of data preparation and positional embeddings.
TREC
Model
Context length
(# of examples)
SPLiCe-no-shuf SPLiCe-shuf
$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
Table 29: Average classification performance on TREC (Li and Roth 2002; Hovy et al. 2001). We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1). SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was the identity function. We use $\pm$ to denote the standard deviation. We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
Appendix K HotPotQA and Qasper Results
We evaluate our models on HotPotQA (Yang et al. 2018) and Qasper (Dasigi et al. 2021) from SCROLLS (Shaham et al. 2022) on both long (see Table 33) and short context (see Table 31 and Table 32)
Task QASPER HotPotQA
Model Context EP SPLiCe Context EP SPLiCe $\Delta$ [conf interv]
[…]
Table 30: We measure question answering over long input using Qasper (Dasigi et al. 2021) (2-shot setting) and HotPotQA (Yang et al. 2018) (10-shot setting). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al. 2021). For HotPotQA, we average results across $7$ sets of in-context examples for $3$B and $7$B models and $2$ for $13$B ones, with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals. We use F1 score for evaluation. Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
QASPER
Model Context Method Result
3B FoT 32K SPLiCe 22.8
2K SPLiCe 18.4
2K EP 18.4
32K ST CHKPT 6.4
2K ST CHKPT 18.4
7B FoT 32K SPLiCe 23.1
2K SPLiCe 18.0
2K EP 18.2
2K ST CHKPT 18.3
7B CL 32K SPLiCe 29.7
2K SPLiCe 18.6
2K EP 18.4
2K ST CHKPT 18.3
13B FoT 32K SPLiCe 32.0
2K SPLiCe 18.9
2K EP 19.1
2K ST CHKPT 19.9
Table 31: Comparison with starting checkpoint on Qasper. We note that the OpenLLaMA models were trained with context length 2K.
HotPotQA
Size Context Method Result
3B FoT 32K SPLiCe 29.6
2K SPLiCe 26.4
2K EP 27.2
2K ST CHKPT 27.3
7B FoT 32K SPLiCe 27.3
2K SPLiCe 21.0
2K EP 19.7
2K ST CHKPT 27.7
7B CL 32K SPLiCe 31.2
2K SPLiCe 24.2
2K EP 24.3
2K ST CHKPT 27.7
13B FoT 2K SPLiCe 36.7
2K SPLiCe 27.6
2K EP 25.6
2K ST CHKPT 32.9
Table 32: Comparison with starting checkpoint on HotPotQA (0-shot - 2k context and 10-shot 20k context). We note that the OpenLLaMA models were trained with context length 2K.
Task Landmark Attention RULER
Model Method Passkey Retrieval Variable Tracking
[…]
EP 68.6 29.8
[…]
@@
We perform an additional evaluation using HumanEval (Chen et al. 2021) and present the results in Table 34. We note that similarly to GSM8K results from Table 3 SPLiCe improves the short context performance of the larger model.
HumanEval
@@
EP 26.8
Table 34: We additionally evaluate our $7$B and $13$B parameter models on HumanEval (Chen et al. 2021). For $7$B models, all methods experience a decrease in performance, but SPLiCe is better at maintaining short context performance. For $13$B, both SPLiCe and EP increase the performance. We note that in our evaluation we do not use special tokens designed for CodeLlama $13$B code completion. Instead, evaluate all models in the same way. Because of the different evaluation pipeline our results for Starting Checkpoint are not the same as the ones presented in (Rozière et al. 2023).
Appendix M $270$M Models Training, Fine-Tuning and Evaluation Details
Training protocol
Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer 2023). Subsequently, we fine-tune using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer 2023) and long context data created using SPLiCe/EP. The amount of pre-training tokens is based on scaling laws from (Hoffmann et al. 2022) and constants for GPT-like models in (Karpathy 2022).
Fine-tuning protocol
We employ the Focused Transformer (FoT) (Tworkowski et al. 2023) for context extension method (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al. 2023; Tworkowski et al. 2023). Loosely inspired by (Ouyang et al. 2022; Rozière et al. 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. We also check how SPLiCe works with other context extension methods in Appendix B.2.
Evaluation
We measure perplexity on held-out portions of the arXiv (Azerbayev, Piotrowski, and Avigad 2022) and StarCoder (Li et al. 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al. 2023; Li et al. 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model. For information regarding the training and evaluation data, see Appendix F.
Appendix N Needle In A Haystack
We utilize the following prompt for evaluating $7$B parameter CL models on Needle In A Haystack (Kamradt 2023). And the default needle that was used by the authors of the benchmark to evaluate models such as Anthropic’s Claude. As a context document, we utilized the “PaulGrahamEssays” option from the Needle In A Haystack evaluator. We use recall of keywords from the ground truth answer as the evaluation metric.
You␣are␣given␣the␣following␣Question␣and␣Context.␣Answer␣the␣Question␣using␣the␣information␣hidden␣inside␣the␣Context.
Question:␣{retrieval_question}
Context:␣{context}
Question:␣{retrieval_question}
Answer:
Appendix O TREC 2K Evaluation
TREC Context 2K (90-shot)
Model ST CHKP EP SPLiCe $\Delta$ [conf interv]
[…]
Table 35: We assess the short-context (2K) performance on TREC (Li and Roth 2002; Hovy et al. 2001). We average across $10$ sets of in-context examples and use $\Delta$ [conf interv] to denote the mean improvement of SPLiCe over the starting checkpoint and its $95\%$ bootstrap confidence intervals.